OpenStudy (anonymous):

solve square root(x^2+4x-12)=3

OpenStudy (anonymous):

sqrt. (x^2+4x-12) = 3 => x^2+4x-12 = 9 => x^2+4x-21 = 0 x=3

OpenStudy (anonymous):

sqrt. (x^2+4x-12) = 3 => x^2+4x-12 = 9 => x^2+4x-21 = 0 x=3